OpenStudy (lilpeter504):

Help me with this rate problem An airplane travels 990 miles in 3 hours with the wind and makes the return trip in 3 hours and 40 minutes. Find the speed of the airplane in still air and the speed of the wind.

OpenStudy (anonymous):

Okay, so for this problem, name two variables, say x and y that represent the speed of the airplane and wind respectively. With two variables, you need two equations. In the first situation, (x+y)3=990

OpenStudy (perl):

nice problem

OpenStudy (lilpeter504):

Two variables like, how many hours and minutes?

OpenStudy (anonymous):

No the speeds of the wind and airplane

OpenStudy (lilpeter504):

Oh yeah, because its asking to find the speed of airplane and wind

OpenStudy (anonymous):

Yes, exactly. For the second equation, the wind and airplane are going opposite directions, so you subtract the speeds. (x-y)3(4/6)=990

OpenStudy (lilpeter504):

So why did you put a 3 after the parenthesis?

OpenStudy (anonymous):

The x+y is a speed so you multiply it by the time of the trip to find the distance of the trip.

OpenStudy (lilpeter504):

So the number 3 is the number for 3 hours

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And in the second equation, 3(4/6) is 3 hours and 40 minutes in hours

OpenStudy (anonymous):

then you just solve for x and y

OpenStudy (lilpeter504):

Wait I don't understand where you gotten 4/6

OpenStudy (lilpeter504):

I'm sorry if I am asking all these questions

OpenStudy (anonymous):

No it's totally fine. I had to express 40 minutes in as hours so I divided it by 60 because there are 60 minutes in an hour

OpenStudy (lilpeter504):

So 40/60 you made it simplified by 10 making it 4/6 correct?

OpenStudy (anonymous):

Yes

OpenStudy (lilpeter504):

Alright could you explain what to do next?

OpenStudy (lilpeter504):

Now that I understand what number stands for what

OpenStudy (anonymous):

Well you have to solve for x and y. To do that, turn the first equation into a y=(990-3x)/3

OpenStudy (anonymous):

Then substitute that into the second equation to get (11/3)x+(11/3)((990-3x)/3)=990

OpenStudy (anonymous):

11/3 is the same as 3(4/6) I just simplified it

OpenStudy (anonymous):

\[\frac{ 11x }{ 3 }+\frac{ 10890-33x }{ 9 }=990\]

OpenStudy (lilpeter504):

Where di you gt -33x

OpenStudy (lilpeter504):

did you get

OpenStudy (anonymous):

Okay wait I sec. I need to rethink this Give me a moment

OpenStudy (lilpeter504):

Because if you were to multiple 11/3 and 990-3x, shouldnt it be 10890/9x

OpenStudy (lilpeter504):

wait 10890/-9x

OpenStudy (anonymous):

its (990-3x)/3

OpenStudy (anonymous):

but wait a moments

OpenStudy (anonymous):

Okay I found an easier way to solve this. Let x equal the speed of the airplane when going with the wind. That means 3*x=990 and x=330

OpenStudy (anonymous):

Let y equal speed of airplane when going against the wind. Thus we have (11/3)*y= 990 and y = 270

OpenStudy (anonymous):

The difference between the two speeds is 60 and that is the difference between when the two speeds when the air is working with and working against, thus when the wind is not there, thus the difference between then the plane is working with the plane and flying in still air is half that distance or 30

OpenStudy (anonymous):

So the speed of the plane in still air 300mph and the speed of the wind is 30mph

OpenStudy (lilpeter504):

Could you reword that above of what you just said before your last response? I didn't quite get what you were saying

OpenStudy (anonymous):

Just think about it as the average of the two speeds

OpenStudy (lilpeter504):

Oh, So you are saying 330 with the wind and then you try to find out how fast the wind is going from the opposite direction which is 30, 330-30 = 300 mph

OpenStudy (lilpeter504):

I still don't quite understand how you got 11/3 by simplifying 3(4/6)

OpenStudy (anonymous):

4/6 =2/3

OpenStudy (anonymous):

3(2/3)=(9+2)/3

OpenStudy (lilpeter504):

Ohh okay, I see where you got it, you just had to make the fraction a mixed number to get 11/3, I keep forgetting my past year math skills

OpenStudy (anonymous):

Cool. So do you understand everything now?

OpenStudy (lilpeter504):

Alright one last question, where did you get 270? Did you divide 11/3 from 990?

OpenStudy (anonymous):

Divide 990 by 11/3 which is the same as multiplying 990 by 3/11

OpenStudy (lilpeter504):

Oh okay, Thank you very much, I understand what you were trying to explain to me

OpenStudy (anonymous):

Awesome. I wish you all the best in all your future math endeavors. May the shwartz be with you.